# Fast and Accurate Charge Transfer Excitations via Nested Aufbau Suppressed Coupled Cluster

### v1

@@ II Acknowledgements
${R_{h\bar{h}}^{p\bar{p}}}^{(n)}$ $=\left(-2\bar{f}_{hp}^{(0)}{t_{h}^{p}}^{(n)}-2\bar{f}^{(0)}_{\bar{h}\bar{p}}{t_{\bar{h}}^{\bar{p}}}^{(n)}\right){t_{h\bar{h}}^{p\bar{p}}}^{(0)}+\left(\bar{f}_{pp}^{(0)}+\bar{f}_{\bar{p}\bar{p}}^{(0)}-\bar{f}_{hh}^{(0)}-\bar{f}_{\bar{h}\bar{h}}^{(0)}-2\bar{f}_{hp}^{(0)}{t_{h}^{p}}^{(0)}-2\bar{f}_{\bar{h}\bar{p}}^{(0)}{t_{\bar{h}}^{\bar{p}}}^{(0)}\right){t_{h\bar{h}}^{p\bar{p}}}^{(n)}+\cdots$ (S9)
@@
S3 Orbital Transformations for Nested Coupled Cluster
@@
The full table of results for the QUEST and charge transfer benchmarks may be found in a .xlsx file available upon request. On rare occasion, an ASCC-based method would experience convergence difficulties, particularly when the ESMF wave function had significant deficiencies as compared to EOM-CCSD. If neither of the two solutions would converge for any of the ASCC-based methods, they were entirely removed for all methods. States where only one of the solutions failed to converge or only a subset of methods failed to converge are highlighted in red. For the states where only one of two solutions would converge, the final reported energy is not the arithmetic mean of the two solutions but just the energy of the only converged solution. Finally, a handful of states would stall near but not quite at our tight convergence criterion. These states, highlighted in blue, are still included due to their energy changing below the level of energetic precision reported in this work.
Three different energetic thresholds for the nesting procedure were tested in this study: 0.0025 eV, 0.005 eV, and 0.01 eV. Ultimately, we decided to report the results of the 0.005 eV threshold in the main text due to its balance of computational efficiency and accuracy. The accuracy for each of the three tested thresholds for the QUEST benchmark are summarized in Figure S2.
@@
S5 Geometries
Geometries for all molecules in the QUEST and CT benchmarks may be found in the Supporting Information from our previous work.[21] The geometries for the large molecule tests may be found below in Angstroms.

### v2

@@ II Acknowledgements
${R_{h\bar{h}}^{p\bar{p}}}^{(n)}$ $=\left(-2\bar{f}_{hp}^{(0)}{t_{h}^{p}}^{(n)}-2\bar{f}^{(0)}_{\bar{h}\bar{p}}{t_{\bar{h}}^{\bar{p}}}^{(n)}\right){t_{h\bar{h}}^{p\bar{p}}}^{(0)}$ (S9)
$+\left(\bar{f}_{pp}^{(0)}+\bar{f}_{\bar{p}\bar{p}}^{(0)}-\bar{f}_{hh}^{(0)}-\bar{f}_{\bar{h}\bar{h}}^{(0)}-2\bar{f}_{hp}^{(0)}{t_{h}^{p}}^{(0)}-2\bar{f}_{\bar{h}\bar{p}}^{(0)}{t_{\bar{h}}^{\bar{p}}}^{(0)}\right){t_{h\bar{h}}^{p\bar{p}}}^{(n)}+\cdots$ (S10)
@@
S3 Orbital Transformations for
Nested Coupled Cluster
@@
The full table of results for the QUEST and charge transfer benchmarks may be found in the included .xlsx file. On rare occasion, an ASCC-based method would experience convergence difficulties, particularly when the ESMF wave function had significant deficiencies as compared to EOM-CCSD. If neither of the two solutions would converge for any of the ASCC-based methods, they were entirely removed for all methods. States where only one of the solutions failed to converge or only a subset of methods failed to converge are highlighted in red. For the states where only one of two solutions would converge, the final reported energy is not the arithmetic mean of the two solutions but just the energy of the only converged solution. Finally, a handful of states would stall near but not quite at our tight convergence criterion. These states, highlighted in blue, are still included due to their energy changing below the level of energetic precision reported in this work.
Three different energetic thresholds for the nesting procedure were tested in this study: 0.0025 eV, 0.005 eV, and 0.01 eV. The accuracy for each of the three tested thresholds for the QUEST benchmark are summarized in Figure S2. Ultimately, we decided to report the results of the 0.005 eV threshold in the main text due to its balance of computational efficiency and accuracy, though we note that the optimal threshold will likely depend on the specifics of the molecular system, excited state character, and basis set. We look forward to clarifying the specifics of this dependence in future work.
@@
S5 State Characterization for Potential Energy Surfaces
In an effort to characterize how accurately each method reproduces the state character of the two CT states in the hydrogen bonding test, we have constructed a rough measure for state characterization. The idea is to quantify the largest component of the state that is not CT, and to compare that between methods, but since we are working with significantly different methods implemented in different codes that print different amounts of state information, a somewhat crude approach is necessary. For the $A^{\prime}$ state, we define the state’s “secondary character” (i.e. the biggest component that is not CT) as (a) the weight of the largest printed single excitation other than the nitrogen-lone-pair-to-$\pi^{*}$ excitation divided by (b) the sum of the weights of all printed single excitations. For the $A^{\prime\prime}$ state, most methods recovered the $\pi^{*}$ orbital from a linear combination of multiple virtuals, meaning that the CT character of the state was constructed using a superposition of multiple single excitations. We therefore define this state’s secondary character as (a) the net weight of all printed single excitations whose holes were not the nitrogen lone pair divided by (b) the net weight of all printed singles. Although the resulting comparison is necessarily somewhat approximate, it does make clear that, in two cases, one for TD-DFT and one for EOM-CCSD, a state was erroneously predicted to contain a substantial amount of non-CT character. As seen in the figure in the main text, both CC3 and nested PLASCC predict that both states have only very small amounts of secondary character.
S6 Geometries
Information regarding geometries for all molecules in the valence and Rydberg QUEST benchmarks as well as the intermolecular CT states and one intramolecular CT state may be found in the Supporting Information from our previous work. [21] Geometries for the remaining intramolecular CT states may be found in the Supporting Information of the work of Loos and coworkers. [34] Geometries for the hydrogen bonding test may be found in the Supporting Information of the work of Clune and coworkers. [52] The remaining geometries for the large molecule tests may be found below in Angstroms.